OpenStudy (anonymous):

Six cans of soda and two bags of potato chips cost $7.08. And eight cans of soda and five bags of potato chips cost $12.45.What is the price per can for the soda, and price per bag of potato chips?

OpenStudy (chaise):

Let s = soda Let c = chips 6s + 2c = 7.08 8s + 5c = 12.45 Multiply the top equation by 4. Multiply the bottom equation times 3. 24s + 8c = 28.32 24s + 15c = 37.35 Take away them from eachother. -7c = -9.03 c = -9.03/7 c = 1.29 Substitute this back into your original equation. If c = 1.29 6s + 2c = 7.08 6s + 2(1.29) = 7.08 6s + 2.58 = 7.08 6s = 4.5 s = 0.75 Chips are $1.29 and cans of soda are $0.75

OpenStudy (chaise):

If you need help with any of this let me know and I would be happy to explain it to you :)